OpenStudy (anonymous):

THE SUM OF EVEN NUMBERS BETWEEN 1 TO 31 please reply quickly

OpenStudy (anonymous):

if you need it fast, get a calculator and add them up

OpenStudy (amistre64):

what does an even number look like in general?

OpenStudy (anonymous):

smooth

OpenStudy (anonymous):

good morning! looks like things have speeded up a bit here

OpenStudy (amistre64):

morning :) and yeah, but i just got an aw snap page nonetheless:/

OpenStudy (amistre64):

even numbers are generally of the form: 2n 2(1) = 2 2(15) = 30 \[\sum_{n=1}^{15}2n\] \[2\sum_{n=1}^{15}n\]

OpenStudy (anonymous):

this is an example of arithmetic progression , so how can i solve it using the formula ?

OpenStudy (amistre64):

which formula?

OpenStudy (amistre64):

2+4+6+8+... common difference of 2 an = a1 + d(n-1) an = 2 + 2(n-1) an = 2(1+n-1) an = 2n ; which gets us to the general form of an even number Sn = n(a1+an)/2 Sn = n(2+2n)/2 Sn = 2n(1+n)/2 Sn = n(1+n)

OpenStudy (anonymous):

thank u , i got the key to get an answer .... can i ask another question that i realy cant understand thank....

OpenStudy (anonymous):

please help me to solve this equation n/2[2a+(n-1)d]=1800 where a=6,d=6,Sn=1800

OpenStudy (amistre64):

\[ \frac n2(2a+(n-1)d)=S_n\] where a=6,d=6,Sn=1800 just replace the letters with the given values \[ \frac n2(2(6)+(n-1)6)=1800\] \[ \frac n2(12+6n-6)=1800\] \[ \frac n2(6+6n)=1800\] \[ \frac n26(1+n)=1800\] \[ 3n(1+n)=1800\]and solve for n